### v4

@@ Abstract
Automated defect detection in transmission lines is essential for maintaining power grid reliability, yet it remains challenging owing to the extreme prevalence of small-scale targets in UAV-acquired imagery. Existing methods rely predominantly on RGB inputs and thus depend on photometric appearance as the primary discriminative cue—a strategy that falters when defects exhibit low chromatic contrast, geometrically ambiguous damage patterns, or partial occlusion by vegetation. Depth imagery offers complementary geometric evidence, but fusing heterogeneous modalities is non-trivial: RGB and depth features differ substantially in noise characteristics and distributional properties, and naïve integration risks propagating modality-specific artifacts into the joint representation. This paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network that follows a principled purify-then-fuse paradigm for RGB-D defect detection. Within each modality branch, a Semantic Recomposition Module projects features through a learned bottleneck with position-wise normalization, suppressing sensor-specific noise—illumination artifacts in RGB and quantization errors in depth—while narrowing the inter-modal distribution gap prior to integration. At the deepest fusion level, a Contextual Semantic Integration Framework establishes global spatial dependencies through partial-channel attention, enabling the network to exploit structural priors such as the regular arrangement of insulator strings for disambiguating small defects from visually similar background elements, without the detail erosion that full-channel attention would impose. Extensive experiments on the TL-RGBD benchmark, where 94.5% of annotated instances qualify as small objects, demonstrate that CMAFNet consistently outperforms state-of-the-art detectors across all evaluation metrics. The full-scale variant achieves 32.2% mAP50 and 12.5% small-object AP, surpassing the strongest baseline by 9.8 and 4.0 percentage points respectively, while the lightweight configuration attains 24.8% mAP50 at 228 FPS with only 4.9M parameters. Ablation studies further reveal a synergistic interaction between the two modules, yielding a combined gain of 13.7% that substantially exceeds their individual contributions, and modality ablation confirms that depth provides complementary geometric cues—particularly for boundary disambiguation and low-contrast defect localization—that RGB appearance alone cannot capture.
@@ 1 Introduction
To address these challenges, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network built on a purify-then-fuse paradigm for RGB-D feature integration. The design is motivated by the observation that multimodal fusion operating on noisy, statistically misaligned representations tends to inherit and amplify their deficiencies rather than compensate for them. CMAFNet therefore interposes a dedicated purification stage before fusion: within each modality branch, a Semantic Recomposition Module (SRM) reconstructs input features through a structured bottleneck that suppresses modality-specific artifacts, such as depth holes and edge bleeding or RGB specular noise, and reduces distributional mismatch between the two feature streams prior to integration. Fusion then operates on these purified representations, and a Contextual Semantic Integration Framework (CSIF) captures long-range structural dependencies that convolutional operators model only at disproportionate cost, a capability important for reasoning about spatial relationships among small defects embedded in densely repetitive structural patterns. To mitigate the detail erosion commonly associated with global attention, CSIF adopts a partial-channel design that balances global context modeling with the fine-grained spatial discrimination required for accurate small-object localization. The resulting architecture is instantiated as a scalable model family, ranging from compact configurations suitable for real-time UAV inference to full-scale variants for server-side post-processing. Experiments on the TL-RGBD benchmark show that CMAFNet consistently outperforms compared methods across evaluation metrics and model scales; for example, the full-scale variant achieves a gain of +2.8 mAP50 and +3.6 mAPS over the strongest baseline under an identical training protocol, while the lightweight variant maintains competitive accuracy at real-time throughput.
The core contributions of this work are threefold. We introduce the purify-then-fuse paradigm for RGB-D feature integration in detection tasks dominated by small objects, instantiated as CMAFNet, a scalable architecture family spanning lightweight to full-scale configurations that achieves consistent improvements across all model scales on TL-RGBD; systematic fusion-strategy ablations confirm that purification before fusion yields measurable gains over direct-fusion alternatives. We propose the Semantic Recomposition Module (SRM), which suppresses modality-specific noise, including depth quantization artifacts and RGB illumination variations, and reduces cross-modal distributional mismatch through a structured reconstruction bottleneck; applied within each modality branch and at fused feature levels, SRM brings heterogeneous RGB and depth representations into more compatible statistical properties, as validated by modality ablations and noise-sensitivity analyses presented in Section 3. We also design the Contextual Semantic Integration Framework (CSIF), a partial-channel global attention mechanism that captures long-range structural dependencies while preserving the spatial detail essential for small-object detection; by restricting global attention to a subset of feature channels and retaining the remainder for local discrimination, CSIF addresses the tension between contextual coverage and spatial precision that limits conventional full-channel attention, as shown through component-level ablation.
@@ 2.1 Overview of Network Architecture
Although integrating depth geometry with RGB appearance holds promise for strengthening defect detection, the effectiveness of such integration depends critically on how the two modalities interact during feature extraction and fusion. Naive combination strategies—whether through early concatenation or late decision fusion—yield limited improvements, as RGB and depth signals differ substantially in statistical properties and noise characteristics (Zhao et al., 2024). Early concatenation forces the network to reconcile heterogeneous distributions from the first layer onward, while late fusion defers cross-modal interaction to the prediction stage and provides little opportunity for mutual enhancement. A more principled strategy is to fuse modality-specific features at intermediate representation levels, where sufficient semantic abstraction exists to tolerate noise while spatial resolution remains adequate for small-object localization. Existing approaches operating at this level, however, typically employ static operations such as element-wise summation or channel concatenation followed by convolution. Such fixed fusion schemes cannot adapt to the dynamic variations inherent in aerial inspection: when RGB features are corrupted by illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these modality-specific corruptions into the shared representation, ultimately degrading rather than enhancing detection performance.
These considerations motivate a fusion paradigm in which each modality stream first undergoes explicit noise suppression and distribution normalization prior to cross-modal integration, and in which the fused representation is subsequently refined through context-aware mechanisms that extend beyond the local receptive field of standard convolutions. The proposed Cross-Modal Alignment and Fusion Network (CMAFNet) instantiates this paradigm by extending the YOLO 11 (Jocher et al., 2023) single-stream detector into a dual-branch architecture for RGB-D input, while retaining all standard components—C3k2, SPPF, the FPN-PAN neck, and the anchor-free decoupled detection head—without modification. Three design elements distinguish CMAFNet from the baseline: (i) a Semantic Recomposition Module (SRM) that performs bottleneck-based feature purification with position-wise distribution normalization, (ii) a Contextual Semantic Integration Framework (CSIF) that introduces attention-driven global context modeling at the deepest fusion level, and (iii) a selective fusion strategy that restricts cross-modal integration to $P_{4}$ and $P_{5}$ while routing $P_{3}$ features exclusively from the RGB branch. Fig. 1 illustrates the complete network topology.
@@
The data flow comprises three stages. In the first stage, the dual-branch backbone encodes RGB and depth inputs independently through to $P_{5}$, with SRM inserted at $P_{3}$ and $P_{4}$ within each branch to suppress modality-specific noise and narrow the inter-modal distribution gap. In the second stage, cross-modal fusion at $P_{4}$ and $P_{5}$ combines the purified branch features through channel concatenation, C3k2 re-encoding, and a second SRM pass; at $P_{5}$, CSIF is additionally interposed between C3k2 and the terminal SRM to capture global context. In the third stage, the detection head receives the fused $P_{4}$$P_{5}$ representations together with the RGB-only $P_{3}$ features through the standard FPN-PAN bidirectional pyramid for multi-scale prediction.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone comprises two encoding branches with identical topology but entirely separate parameters, one processing RGB input and the other processing depth input. All convolutional layers, C3k2 blocks, and the terminal SPPF module adopt the default YOLO 11 (Jocher et al., 2023) implementation without modification.
The RGB branch receives 3-channel images, while the depth branch receives single-channel depth maps. Depth maps in the TL-RGBD dataset are spatially registered with their RGB counterparts through the factory calibration of the onboard RGB-D sensor (Intel RealSense D435i), and no additional geometric alignment is performed during preprocessing. Missing depth values and edge artifacts inherent to the sensor are not explicitly inpainted; robustness to these corruptions is instead delegated to the SRM purification and the selective fusion strategy described in Section 2.1.2.
For the nano-scale configuration (width_multiple =0.25=0.25, depth_multiple =0.50=0.50, input $640\times 640$), the feature hierarchy spans five pyramid levels: $P_{1}$ ($320{\times}320$, C=16C{=}16), $P_{2}$ ($160{\times}160$, C=32C{=}32), $P_{3}$ ($80{\times}80$, C=64C{=}64), $P_{4}$ ($40{\times}40$, C=128C{=}128), and $P_{5}$ ($20{\times}20$, C=256C{=}256), with strides of 2, 4, 8, 16, and 32, respectively. The two branches produce identical feature dimensions at every level; the sole asymmetry is that the output of $P_{3}$ in the depth branch is not propagated to the detection head.
Within each branch, an SRM instance (K=128K{=}128, $\alpha{=}0.8$; see Section 2.2) is inserted immediately after the C3k2 block at both $P_{3}$ and $P_{4}$. These scales strike a favorable balance: semantic content is sufficiently rich for effective refinement, yet spatial resolution remains fine enough to preserve small-defect detail. Applying SRM before fusion narrows the inter-modal distribution gap at the source, thereby providing cleaner inputs for the subsequent concatenation-based integration.
The rationale for adopting parallel branches, rather than early fusion, lies in the distinct noise profiles of the two modalities. RGB imagery exhibits illumination variation and specular reflections, whereas depth maps suffer from edge bleeding, missing values at reflective surfaces, and quantization noise. Merging these signals before individual processing would entangle their noise distributions and complicate subsequent alignment. Independent encoding through $P_{5}$ allows each branch to develop modality-appropriate representations, deferring cross-modal integration to $P_{4}$ and $P_{5}$, where greater semantic abstraction renders noise-induced misalignment less problematic.
2.1.2 Fusion Layers: Selective Cross-Modal Integration at $P_{4}$ and $P_{5}$
@@
Two observations motivate the exclusion of depth from $P_{3}$. First, depth features at this resolution carry the highest level of sensor noise and spatial misalignment; injecting them into the feature map most critical for small-defect localization would introduce more interference than useful geometric cues. Second, the primary contribution of depth—surface discontinuities and inter-component spatial layout—is adequately captured at $P_{4}$ and $P_{5}$, where feature granularity better matches the effective resolution of the sensor.
@@ 2.1.3 Detection Head
Each scale terminates in a decoupled head with separate classification and regression branches. Bounding boxes are predicted through distribution focal loss (DFL), which regresses discretized distributions over left-top-right-bottom offsets relative to anchor points. Positive samples are assigned by the Task-Aligned Assigner, which scores candidates jointly on the basis of classification confidence and localization quality. The training loss combines binary cross-entropy for classification, CIoU for box regression, and DFL for distribution regression, all with default YOLO 11 weights. The proposed contributions are thus confined entirely to the feature extraction and fusion stages; the detection head and training procedure remain identical to those of the baseline.
@@ 2.2 Semantic Recomposition Module
SRM is designed to suppress modality-specific noise without eroding the fine-grained details on which small-defect localization depends. In transmission-line imagery, background textures such as conductor strand patterns and fitting reflections, structural edges at conductor–sky boundaries, and depth sensor artifacts including holes, edge bleeding, and quantization noise all produce strong feature responses that inflate false-positive rates at the scales most relevant to small targets. Channel and spatial attention mechanisms can reweight existing activations but cannot impose structural constraints on what constitutes valid content. SRM adopts a fundamentally different strategy: it compresses features into a learned low-dimensional subspace, refines them within that subspace, and reinjects the result through controlled residual mixing.
@@
The reduction from $C$ to $K$ introduces an information bottleneck that retains only components admitting a stable low-dimensional representation while discarding incidental variation.
@@
No activation function follows this convolution; maintaining linearity in the refinement stage prevents nonlinear coupling of noise across latent channels. The $5\times 5$ kernel is sufficiently large to capture local relationships among adjacent structural elements while preserving parameter efficiency through the depthwise formulation.
@@
where $\epsilon=10^{-5}$. Unlike batch or layer normalization, which pool statistics over spatial or batch dimensions, position-wise normalization computes statistics at each $(h,w)$ location independently. Transmission-line images exhibit pronounced spatial heterogeneity—sky regions, metallic towers, and insulator assemblies produce markedly different activation patterns—and sharing statistics across these regions would suppress informative variation. A consequence of this design is that absolute activation magnitudes are erased at each position, leaving only the relative distribution across latent channels. This property reduces sensitivity to the systematic scale discrepancy between RGB and depth activations at corresponding locations, yet it also discards magnitude information that may be relevant to confidence estimation. The residual mixing in the final stage partially compensates for this loss by preserving a fraction of the original features.
@@
where $\alpha=0.8$ is a fixed scalar. Assigning a large weight to the refined branch maximizes noise suppression, while the $(1{-}\alpha)$ residual preserves fine-grained spatial detail that the bottleneck cannot fully reconstruct, thereby striking a balance between purification strength and localization fidelity.
All six instances share the same architecture but serve distinct functional roles determined by the choice of $K$. At the branch level (K=128K{=}128), the compact bottleneck filters modality-private noise—illumination artifacts in RGB and sensor holes together with quantization errors in depth—by compressing features before normalization, yielding per-modality representations with narrower inter-modal distribution gaps. At the fusion level (K=256K{=}256), the expanded bottleneck operates on concatenated cross-modal features; position-wise normalization then acts on the joint RGB-depth latent space, attenuating channels whose activations are inconsistent across modalities while reinforcing those that exhibit coherent patterns. In both cases, $K$ is set to approximately half the input channel dimension, balancing compression strength against representational capacity.

### v5

@@ Abstract
Automated defect detection in transmission lines is critical for power grid reliability but remains challenging because small-scale targets dominate UAV-acquired imagery. Most methods depend on RGB inputs alone, treating photometric appearance as the primary cue, a strategy that fails when defects show low chromatic contrast, ambiguous damage geometry, or partial occlusion by vegetation. Depth imagery offers complementary geometric evidence, yet fusing it with RGB is non-trivial: the two modalities differ markedly in noise characteristics and distributional properties, so naïve combination can propagate modality-specific artifacts into the joint representation. We propose CMAFNet, a Cross-Modal Alignment and Fusion Network built on a purify-then-fuse paradigm for RGB-D defect detection. A Semantic Recomposition Module within each branch projects features through a learned bottleneck with position-wise normalization, suppressing sensor-specific noise while narrowing the inter-modal distribution gap. At the deepest fusion stage, a Contextual Semantic Integration Framework models global spatial dependencies via partial-channel attention, enabling the network to use structural priors such as regular insulator-string arrangement to distinguish small defects from similar background elements without the detail loss of full-channel attention. On the TL-RGBD benchmark, where 94.5% of annotated instances are small objects, CMAFNet surpasses state-of-the-art detectors on all metrics. The full-scale model achieves 32.2% mAP50 and 12.5% small-object AP, exceeding the best baseline by 9.8 and 4.0 points; the lightweight variant reaches 24.8% mAP50 at 228 FPS with 4.9M parameters. Ablation studies show the two modules interact synergistically with a combined gain of 13.7% that exceeds their individual contributions, and confirm that depth provides geometric cues for boundary disambiguation and low-contrast localization that RGB alone cannot supply.
@@ 1 Introduction
To address these challenges, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network that follows a purify-then-fuse paradigm for RGB-D feature integration. The core motivation is straightforward: fusing noisy, statistically misaligned representations tends to propagate their deficiencies rather than compensate for them. CMAFNet therefore introduces a dedicated purification stage before fusion. Within each modality branch, a Semantic Recomposition Module (SRM) reconstructs features through a structured bottleneck that suppresses modality-specific artifacts such as depth holes, edge bleeding, and RGB specular noise, while reducing distributional mismatch between the two streams. A Contextual Semantic Integration Framework (CSIF) then fuses these purified representations by capturing long-range structural dependencies that convolutional operators can only model at disproportionate cost, which is particularly relevant for reasoning about spatial relationships among small defects in densely repetitive patterns. To mitigate the detail erosion commonly associated with global attention, CSIF adopts a partial-channel design that balances contextual modeling with the fine-grained spatial discrimination needed for small-object localization. The resulting architecture is instantiated as a scalable model family, from compact configurations for real-time UAV inference to full-scale variants for server-side post-processing.
The core contributions of this work are threefold:
  • 1.
    We introduce the purify-then-fuse paradigm for RGB-D feature integration in small-object detection and instantiate it as CMAFNet, a scalable architecture family spanning lightweight to full-scale configurations. CMAFNet achieves consistent improvements across all model scales on TL-RGBD, and systematic ablations confirm that purification before fusion yields measurable gains over direct-fusion alternatives.
  • 2.
    We propose the Semantic Recomposition Module (SRM), which suppresses modality-specific noise such as depth quantization artifacts and RGB illumination variations through a structured reconstruction bottleneck. Applied within each modality branch and at fused feature levels, SRM brings heterogeneous RGB and depth representations into more compatible statistical distributions, as validated by modality ablations and noise-sensitivity analyses in Section 3.
  • 3.
    We design the Contextual Semantic Integration Framework (CSIF), a partial-channel global attention mechanism that captures long-range structural dependencies while preserving spatial detail for small-object detection. By restricting global attention to a channel subset and retaining the remainder for local discrimination, CSIF addresses the tension between contextual coverage and spatial precision that limits conventional full-channel attention, as confirmed by component-level ablation.
@@ 2.1 Overview of Network Architecture
Combining depth geometry with RGB appearance can strengthen defect detection, but the gains depend heavily on how the two modalities interact during feature extraction and fusion. Naive strategies, whether early concatenation or late decision fusion, offer limited improvement because RGB and depth signals differ substantially in their statistical properties and noise characteristics (Zhao et al., 2024). Early concatenation forces the network to reconcile heterogeneous distributions from the first layer onward, while late fusion defers cross-modal interaction to the prediction stage, leaving little room for the modalities to inform each other. A more effective strategy is to fuse modality-specific features at intermediate levels, where semantic abstraction is sufficient to tolerate noise yet spatial resolution still supports small-object localization. Prior methods that operate at this level, however, typically rely on static operations such as element-wise summation or channel concatenation followed by convolution. These fixed schemes cannot accommodate the dynamic variations common in aerial inspection: when RGB features are corrupted by illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these corruptions into the shared representation, degrading detection performance rather than improving it.
These limitations call for a fusion design in which each modality stream first undergoes explicit noise suppression and distribution normalization before cross-modal integration, and in which the fused representation is then refined through context-aware mechanisms that reach beyond the local receptive field of standard convolutions. The proposed Cross-Modal Alignment and Fusion Network (CMAFNet) realizes this idea by extending the YOLO 11 (Jocher et al., 2023) single-stream detector into a dual-branch architecture for RGB-D input, while keeping all standard components, including C3k2, SPPF, the FPN-PAN neck, and the anchor-free decoupled detection head, unchanged. Three elements distinguish CMAFNet from the baseline: a Semantic Recomposition Module (SRM) that performs bottleneck-based feature purification with position-wise distribution normalization, a Contextual Semantic Integration Framework (CSIF) that introduces attention-driven global context modeling at the deepest fusion level, and a selective fusion strategy that restricts cross-modal integration to $P_{4}$ and $P_{5}$ while routing $P_{3}$ features exclusively from the RGB branch. Fig. 1 shows the complete network topology.
@@
The data flow proceeds in three stages. In the first, the dual-branch backbone encodes RGB and depth inputs independently through to $P_{5}$, with SRM inserted at $P_{3}$ and $P_{4}$ within each branch to suppress modality-specific noise and narrow the inter-modal distribution gap. In the second, cross-modal fusion at $P_{4}$ and $P_{5}$ combines the purified branch features through channel concatenation, C3k2 re-encoding, and a second SRM pass; at $P_{5}$, CSIF is additionally placed between C3k2 and the terminal SRM to capture global context. In the third, the detection head receives the fused $P_{4}$$P_{5}$ representations together with the RGB-only $P_{3}$ features through the standard FPN-PAN bidirectional pyramid for multi-scale prediction.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone consists of two encoding branches with identical topology but entirely separate parameters, one processing RGB input and the other processing depth input. All convolutional layers, C3k2 blocks, and the terminal SPPF module follow the default YOLO 11 (Jocher et al., 2023) implementation without modification.
The RGB branch receives 3-channel images, while the depth branch receives single-channel depth maps. Depth maps in the TL-RGBD dataset are spatially registered with their RGB counterparts through the factory calibration of the onboard RGB-D sensor (Intel RealSense D435i), and no additional geometric alignment is applied during preprocessing. Missing depth values and edge artifacts inherent to the sensor are not explicitly inpainted; robustness to these corruptions is instead delegated to SRM purification and the selective fusion strategy described in Section 2.1.2.
For the nano-scale configuration (width_multiple =0.25=0.25, depth_multiple =0.50=0.50, input $640\times 640$), the feature hierarchy spans five pyramid levels: $P_{1}$ ($320{\times}320$, C=16C{=}16), $P_{2}$ ($160{\times}160$, C=32C{=}32), $P_{3}$ ($80{\times}80$, C=64C{=}64), $P_{4}$ ($40{\times}40$, C=128C{=}128), and $P_{5}$ ($20{\times}20$, C=256C{=}256), with strides of 2, 4, 8, 16, and 32, respectively. The two branches produce identical feature dimensions at every level; the only asymmetry is that $P_{3}$ output from the depth branch is not propagated to the detection head.
Within each branch, an SRM instance (K=128K{=}128, $\alpha{=}0.8$; see Section 2.2) is inserted immediately after the C3k2 block at both $P_{3}$ and $P_{4}$. These scales offer a good balance: semantic content is rich enough for effective refinement, yet spatial resolution remains fine enough to preserve small-defect detail. Applying SRM before fusion narrows the inter-modal distribution gap at the source, providing cleaner inputs for the subsequent concatenation-based integration.
The rationale for parallel branches, rather than early fusion, lies in the distinct noise profiles of the two modalities. RGB imagery exhibits illumination variation and specular reflections, whereas depth maps suffer from edge bleeding, missing values at reflective surfaces, and quantization noise. Merging these signals before individual processing would entangle their noise distributions and complicate subsequent alignment. Independent encoding through $P_{5}$ allows each branch to develop modality-appropriate representations, deferring cross-modal integration to $P_{4}$ and $P_{5}$, where greater semantic abstraction makes noise-induced misalignment less problematic.
2.1.2 Fusion Layers
@@
Two observations motivate excluding depth from $P_{3}$. Depth features at this resolution carry the highest level of sensor noise and spatial misalignment; injecting them into the feature map most critical for small-defect localization would introduce more interference than useful geometric cues. Meanwhile, the primary contribution of depth, namely surface discontinuities and inter-component spatial layout, is adequately captured at $P_{4}$ and $P_{5}$, where feature granularity better matches the effective resolution of the sensor.
@@ 2.1.3 Detection Head
Each scale terminates in a decoupled head with separate classification and regression branches. Bounding boxes are predicted through distribution focal loss (DFL), which regresses discretized distributions over left-top-right-bottom offsets relative to anchor points. Positive samples are assigned by the Task-Aligned Assigner, which scores candidates jointly on classification confidence and localization quality. The training loss combines binary cross-entropy for classification, CIoU for box regression, and DFL for distribution regression, all with default YOLO 11 weights. The proposed contributions are thus confined entirely to the feature extraction and fusion stages; the detection head and training procedure remain identical to those of the baseline.
@@ 2.2 Semantic Recomposition Module
SRM suppresses modality-specific noise without eroding the fine-grained details on which small-defect localization depends. In transmission-line imagery, background textures such as conductor strand patterns and fitting reflections, structural edges at conductor–sky boundaries, and depth sensor artifacts including holes, edge bleeding, and quantization noise all produce strong feature responses that inflate false-positive rates at the scales most relevant to small targets. Channel and spatial attention mechanisms can reweight existing activations but cannot impose structural constraints on what constitutes valid content. SRM takes a different approach: it compresses features into a learned low-dimensional subspace, refines them there, and reinjects the result through controlled residual mixing.
@@
The reduction from $C$ to $K$ creates an information bottleneck that retains only components with a stable low-dimensional representation while discarding incidental variation.
@@
No activation function follows this convolution; keeping the refinement stage linear prevents nonlinear coupling of noise across latent channels. The $5\times 5$ kernel is large enough to capture local relationships among adjacent structural elements while remaining parameter-efficient thanks to the depthwise formulation.
@@
where $\epsilon=10^{-5}$. Unlike batch or layer normalization, which pool statistics over spatial or batch dimensions, position-wise normalization computes statistics at each $(h,w)$ location independently. Transmission-line images exhibit strong spatial heterogeneity: sky regions, metallic towers, and insulator assemblies produce markedly different activation patterns, and sharing statistics across these regions would suppress informative variation. A side effect of this design is that absolute activation magnitudes are lost at each position, leaving only the relative distribution across latent channels. This reduces sensitivity to the systematic scale gap between RGB and depth activations at corresponding locations, but also discards magnitude information that may matter for confidence estimation. The residual mixing in the final stage partly compensates for this loss by retaining a fraction of the original features.
@@
where $\alpha=0.8$ is a fixed scalar. Giving most of the weight to the refined branch maximizes noise suppression, while the $(1{-}\alpha)$ residual preserves fine-grained spatial detail that the bottleneck cannot fully reconstruct, balancing purification strength against localization fidelity.
All six instances share the same architecture but play distinct roles depending on $K$. At the branch level (K=128K{=}128), the compact bottleneck filters modality-private noise, such as illumination artifacts in RGB and sensor holes together with quantization errors in depth, by compressing features before normalization. This yields per-modality representations with narrower inter-modal distribution gaps. At the fusion level (K=256K{=}256), the wider bottleneck operates on concatenated cross-modal features; position-wise normalization then acts on the joint RGB-depth latent space, attenuating channels whose activations are inconsistent across modalities while reinforcing those that show coherent patterns. In both cases, $K$ is set to roughly half the input channel dimension, trading off compression strength against representational capacity.